# latexText page

### v1

@@
${}^{1}$raduate School of Economics
@@
Capital allocation is a procedure for quantifying the contribution of each source of risk to aggregated risk. The gradient allocation rule, also known as the Euler principle, is a prevalent rule of capital allocation under which the allocated capital captures the diversification benefit of the marginal risk as a component of overall risk. This research concentrates on Expected Shortfall (ES) as a regulatory standard and focuses on the gradient allocations of ES, also called ES contributions. We achieve the comprehensive treatment of backtesting the tuple of ES contributions in the framework of the traditional and comparative backtests based on the concepts of joint identifiability and multi-objective elicitability. For robust forecast evaluation against the choice of scoring function, we further develop Murphy diagrams for ES contributions as graphical tools to check whether one forecast dominates another under a class of scoring functions. Finally, leveraging the recent concept of multi-objective elicitability, we propose a novel semiparametric model for forecasting dynamic ES contributions based on a compositional regression model. In an empirical analysis of stock returns we evaluate and compare a variety of models for forecasting dynamic ES contributions and demonstrate the outstanding performance of the proposed model.
Keywords: Capital allocation; Compositional data analysis; Elicitability; Euler principle; Expected shortfall.
@@ 1 Introduction
Capital allocation is a second stage for quantifying the contribution of each source of risk to total risk. Reflecting the diversification benefit, the allocated capital to each component is calculated so that the sum of the allocated capital over all components equals the total capital. Among various allocation rules proposed in the literature, the Euler principle is a prevalent allocation rule justified from various perspectives, such as RORAC (return on risk-adjusted capital) compatibility (Tasche 1999), cooperative game theory (Denault 2001), and axiomatic foundation of capital allocations (Kalkbrener 2005); see also Tasche (2008) and references therein. Together with the growing importance of ES in recent banking regulations, this paper focuses on the situation of capital allocation when total risk is measured by ES and allocated capital is calculated under the Euler principle; see Section 2.1 for details. We call the allocated capital calculated in this setting ES contribution.
@@
In contrast to the backtesting of risk measures, studies on backtesting of dynamic capital allocation are quite limited in the literature. Bielecki et al. (2020) propose a backtesting method of ES contributions based on the concept called fairness. This method is essentially treated in the more general framework of calibration and traditional backtests introduced in Nolde and Ziegel (2017). Despite the usefulness of such traditional backtests for model validation, Nolde and Ziegel (2017) also point out several issues of such tests particularly in terms of model comparison and banking regulation. To this end, our first aim in this paper is to achieve the comprehensive treatment of backtesting ES contributions in the framework of traditional and comparative backtests (Nolde and Ziegel 2017) based on the concepts of joint identifiability and multi-objective elicitability (Fissler and Hoga 2023).
@@
Although the results shown in Fissler and Hoga (2023) are on systemic risk measures such as CoVaR, CoES and MES, in Section 2.2 we translate their results under the setting of capital allocation. Summarizing their results in our setting, ES contributions are jointly identifiable with total VaR (i.e., VaR of the aggregated loss); moreover, ES contributions themselves are not elicitable, but are multi-objective elicitable, when combined with total VaR and with respect to the lexicographical order. The last statement means more precisely that the tuple of true ES contributions and total VaR is elicited as a unique minimizer of an expectation of some $\mathbb{R}^{2}$-valued scoring function with respect to the order on $\mathbb{R}^{2}$ such that the forecast evaluation of VaR is prioritized, and that of ES contributions is conducted secondarily. Therefore, in the framework of a comparative backtest based on multi-objective elicitability it is necessary to forecast ES contributions together with total VaR, and two forecasts of ES contributions are comparable only when the corresponding forecasts of total VaR are equally accurate. Section 2.3 summarizes the framework of backtests of dynamic ES contributions. We also discuss order sensitivity of scoring functions of ES contributions, which addresses comparability of two misspecified forecasts.
@@
When conducting a comparative backtest in practice, it may not always be clear which scoring function to use. In addition, the resulting forecast ranking can be sensitive to the choice of scoring functions (Patton 2020). To overcome these problems, a diagnostic tool called the Murphy diagram is explored in  Ehm et al. (2016) and Ziegel et al. (2020) for robust forecast evaluation against the choice of scoring function. Based on a mixture representation of the scoring function, the Murphy diagram is particularly useful for checking whether one forecast dominates another under a class of scoring functions.
@@
As a second aim of this paper, we complement the framework of backtesting ES contributions by developing their Murphy diagrams. They serve as a visual, intuitive way to compare competing forecasting models, thereby making it easier to digest and make informed decisions based on their performance. Section 3 introduces Murphy diagrams of ES contributions and summarizes their properties.
@@
Our last aim of the paper is to leverage the concept of multi-objective elicitability and introduce a novel semiparametric model of dynamic ES contributions. Section 4 proposes such a model, which combines the joint semiparametric estimation of the pair of total VaR and total ES with the compositional regression for modeling the dynamics of the proportions of total ES to each component of risk. Based on multi-objective elicitability of ES contributions, forecast accuracy of total VaR is first evaluated prior to that of ES contributions. In addition, risk allocation is conceptually a second stage in the risk management process after quantifying overall risk. This theoretical and practical two-stage procedure motivates us to first model the dynamics of the pair of total VaR and total ES and then estimate ES contributions through the proportion of each source of risk to total ES. Since the component-wise sum of the vector of proportions must equal $1$, we model its dynamics by the compositional regression, which is a multiple regression for such compositional data.
@@
Note that Boonen et al. (2019) also apply compositional data analysis to the problem of capital allocation. They first estimate allocated capital from the losses under the assumption of normality and then fit and evaluate compositional regression models by regarding the set of normalized allocated capital as compositional data. One potential challenge of this procedure is that statistical error and model uncertainty reside not only in the compositional regression model, but also in the estimated allocated capital. To overcome this issue, we fit and evaluate compositional regression models based on a scoring function of ES contributions. Multi-objective elicitability of ES contributions allows us to quantify the accuracy of compositional regression models based on realized losses and not on estimated allocated capital. The major advantage of our proposed semiparametric model is in the complete separation of modeling ES contributions from that of total VaR and total ES. This feature enables us to concentrate on modeling ES contributions based on the best existing model for the pair of total VaR and total ES and to assess the forecast accuracy of ES contributions among the proposed compositional regression models and an existing one by equalizing their total VaR and total ES.
@@
For the remaining sections and appendices, Section 5 conducts an empirical analysis of a portfolio of stock returns and demonstrates superior performance of the proposed model compared with other models including those with conditional heteroskedastic volatilities, hysteretic effects, and time-varying correlations. Section 6 discusses potential directions for future research. We defer all technical results to Section S1, where references starting with the prefix “S” refer to the supplementary material. We also present details and additional results of our empirical analysis in Section S2.
@@ 2.1 Gradient allocations of expected shortfall
Throughout the paper we treat all vectors as row vectors and add the superscript $\top$ to indicate a column vector or the transpose of a matrix. In addition, we fix an atomless probability space $(\Omega,{\cal A},\mathbb{P})$, where all random objects are defined. For $p\in\{0\}\cup[1,\infty)$ and $d\in\mathbb{N}$, let ${\cal L}^{p}(\mathbb{R}^{d})$ be the set of all $\mathbb{R}^{d}$-value random vectors on $(\Omega,{\cal F},\mathbb{P})$ whose components have a finite $p$th moment. Let $F_{\bm{X}}$ be the joint cumulative distribution function (cdf) of $\bm{X}\in{\cal L}^{0}(\mathbb{R}^{d})$. We also define the collection ${\cal F}^{p}(\mathbb{R}^{d})=\{F_{\bm{X}}:\bm{X}\in{\cal L}^{p}(\mathbb{R}^{d})\}$ for $p\in\{0\}\cup[1,\infty)$ and $d\in\mathbb{N}$. Denote by ${\cal F}_{\text{c}}^{p}(\mathbb{R}^{d})$ the class of cdfs $F$ with a strictly positive Lebesgue density for every $\bm{x}\in\mathbb{R}^{d}$ such that $F(\bm{x})\in(0,1)$. Analogously, ${\cal L}_{\text{c}}^{p}(\mathbb{R}^{d})$ denotes the set of $\mathbb{R}^{d}$-valued random vectors whose cdf belongs to ${\cal F}_{\text{c}}^{p}(\mathbb{R}^{d})$. For $X\in{\cal L}^{0}(\mathbb{R})$, Value-at-Risk (VaR) with a confidence level $\alpha\in(0,1)$ is $\operatorname{VaR}_{\alpha}(X)=\inf\{x\in\mathbb{R}:F_{X}(x)\geq\alpha\}$. Moreover, for $X\in{\cal L}^{1}(\mathbb{R})$, the expected shortfall (ES) with a confidence level $\alpha\in(0,1)$ is $\operatorname{ES}_{\alpha}(X)=\frac{1}{1-\alpha}\int_{\alpha}^{1}\operatorname{VaR}_{\beta}(X){\rm d}\beta$, which coincides with $\operatorname{ES}_{\alpha}(X)=\mathbb{E}[X|X\geq\operatorname{VaR}_{\alpha}(X)]$ if $X\in\mathcal{L}_{\text{c}}^{1}(\mathbb{R})$. The confidence level $\alpha$ is typically chosen to be close to $1$, such as $0.975$ and $0.99$. Let $\bm{X}=(X_{1},\dots,X_{d})\in{\cal L}^{0}(\mathbb{R}^{d})$ be a random vector standing for the collection of losses to a portfolio or a financial institution. Moreover, let $S=X_{1}+\cdots+X_{d}$ be the total loss. Our sign convention is that positive values are losses and negative values are profits.
Under the prevalent Euler principle, the contribution of each component $X_{j}$, $j=1\dots,d$, to the risk of the total loss $\operatorname{ES}_{\alpha}(S)$ is determined by the gradient $\frac{\partial}{\partial\lambda_{j}}\operatorname{ES}_{\alpha}(\bm{\lambda}^{\top}\bm{X})|_{\bm{\lambda}=\mathbf{1}_{d}}$ provided that the partial derivative exists, where $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{d})\in\mathbb{R}^{d}$ and $\mathbf{1}_{d}=(1,\dots,1)\in\mathbb{R}^{d}$. Given certain smoothness conditions (Tasche 1999), this derivative leads to:
@@
which we call the ES contribution for the $j$th risk. Note that the ES contribution itself is well-defined for $(X_{j},S)\in{\cal L}^{1}(\mathbb{R}^{2})$ without smoothness assumptions. If $S\in\mathcal{L}_{\text{c}}^{1}(\mathbb{R})$, then the following full allocation property holds:
@@ 2.2 Multi-objective elicitability and joint identifiability of ES contributions
We introduce several properties required for backtesting ES contributions. Let ${\cal F}\subseteq{\cal F}^{0}(\mathbb{R}^{d})$, $k,\,m\in\mathbb{N}$, and $A\subseteq\mathbb{R}^{k}$. A function $\mathbf{S}:A\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is called $\mathcal{F}$-integrable if, for every $F\in\mathcal{F}$ and $\bm{r}\in A$, every component of the function $\bm{x}\mapsto\mathbf{S}(\bm{r},\bm{x})$ is integrable with respect to $F$. A functional $\bm{\varrho}:{\cal F}\rightarrow A$ is called multi-objective elicitable on ${\cal F}$ with respect to a total order $\preceq$ on $\mathbb{R}^{m}$ if there exists an $\mathcal{F}$-integrable function $\mathbf{S}:A\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ such that, for every $F\in{\cal F}$, $\bm{\varrho}(F)$ is the unique minimizer of $\bm{r}\mapsto\mathbb{E}[\mathbf{S}(\bm{r},\bm{X})]$, $\bm{X}\sim F$, over $A$. The function $\mathbf{S}$ is called a (strictly $\mathcal{F}$-consistent multi-objective) scoring function for $\bm{\varrho}$. We simply call $\bm{\varrho}$ elicitable when a scoring function can be taken with $m=1$.
@@
For $j\in\{1,\dots,d\}$, the $j$th ES contribution (1) coincides with the marginal expected shortfall (MES) of $(S,X_{j})$ considered in Fissler and Hoga (2023) provided that $(S,X_{j})\in\mathcal{F}_{\text{c}}^{1}(\mathbb{R}^{2})$. This observation immediately yields the following results on the multi-objective elicitability and joint identifiability of ES contributions shown in Theorem 4.2 and Theorem S.3.1 of  Fissler and Hoga (2023), respectively. In the following, the lexicographic order $\leq_{\text{lex}}$ is adopted as a total order on $\mathbb{R}^{2}$. For every $(a_{1},b_{1}),(a_{2},b_{2})\in\mathbb{R}^{2}$, we write $(a_{1},b_{1})\leq_{\text{lex}}(a_{2},b_{2})$ if $a_{1}<a_{2}$ or if $(a_{1}=a_{2}\text{ and }b_{1}\leq b_{2})$. Moreover, we define:
@@
    $\mathbf{S}_{j}((m_{j},v),\bm{x})$ $=(\operatorname{S}^{\text{VaR}}(v,s),\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x})),$
@@
    where $s=\sum_{i=1}^{d}x_{i}$, $h:\mathbb{R}\rightarrow\mathbb{R}$ is a strictly increasing function, and $\phi_{j}:\mathbb{R}\rightarrow\mathbb{R}$ is a strictly convex function with subgradient $\phi_{j}^{\prime}$ such that $\mathbf{S}_{j}$ is $\mathcal{F}$-integrable.
@@
    $\mathbf{S}((\bm{m},v),\bm{x})=\left(\operatorname{S}^{\text{VaR}}(v,s),\sum_{j=1}^{d}\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x})\right),$
@@
    $\mathbf{V}_{j}((m_{j},v),\bm{x})$ $=(\operatorname{V}^{\text{VaR}}(v,s),\operatorname{V}_{j}^{\text{ESC}}((m_{j},v),\bm{x})),$
@@
    $\mathbf{V}((\bm{m},v),\bm{x})$ $=(\operatorname{V}^{\text{VaR}}(v,s),\operatorname{V}_{1}^{\text{ESC}}((m_{1},v),\bm{x}),\dots,\operatorname{V}_{d}^{\text{ESC}}((m_{d},v),\bm{x})),$
@@
As a concrete example of the scoring functions in Proposition 1, the choice $h(s)=s$ in $\operatorname{S}^{\text{VaR}}$ yields the well-known pinball loss $\operatorname{S}^{\text{VaR}}(v,s)=(\mathbbm{1}{\{s\leq v\}}-\alpha)(v-s)$, and $\phi_{j}(x)=x^{2}$ in $\operatorname{S}_{j}^{\text{ESC}}$ leads to the square loss $\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}\{s>v\}(x_{j}-m_{j})^{2}$; see Nolde and Ziegel (2017) and Fissler and Hoga (2023) for more examples.

### v2

@@
1Graduate School of Economics
@@
Capital allocation is a procedure for quantifying the contribution of each source of risk to aggregated risk. The gradient allocation rule, also known as the Euler principle, is a prevalent rule of capital allocation under which the allocated capital captures the diversification benefit of the marginal risk as a component of overall risk. This research concentrates on Expected Shortfall (ES) as a regulatory standard and focuses on the gradient allocations of ES, also called ES contributions (ESCs). We present the comprehensive treatment of backtesting the tuple of ESCs in the framework of the traditional and comparative backtests based on the concepts of joint identifiability and multi-objective elicitability. For robust forecast evaluation against the choice of scoring function, we also extend the Murphy diagram, a graphical tool to check whether one forecast dominates another under a class of scoring functions, to the case of ESCs. Finally, leveraging the recent concept of multi-objective elicitability, we propose a novel semiparametric model for forecasting dynamic ESCs based on a compositional regression model. In an empirical analysis of stock returns we evaluate and compare a variety of models for forecasting dynamic ESCs and demonstrate the outstanding performance of the proposed model.
JEL classification: C02, C51, C52, C53, G22.
Keywords: Capital allocation; Compositional data analysis; Dynamic risk management; Elicitability; Euler principle; Expected shortfall.
@@ 1 Introduction
Capital allocation is a second stage for quantifying the contribution of each source of risk to total risk. Reflecting the diversification benefit, the allocated capital to each component is calculated so that the sum of the allocated capital over all components equals the total capital. Among various allocation rules proposed in the literature, the Euler principle is a prevalent allocation rule justified from various perspectives, such as RORAC (return on risk-adjusted capital) compatibility (Tasche 1999), cooperative game theory (Denault 2001), and axiomatic foundation of capital allocations (Kalkbrener 2005); see also Tasche (2008) and references therein. Together with the growing importance of ES in recent banking regulations, this paper focuses on the situation of capital allocation when total risk is measured by ES and allocated capital is calculated under the Euler principle; see Section 2.1 for details. We call the allocated capital calculated in this setting ES contribution (ESC).
@@
In contrast to the case of risk measures, studies on backtesting of dynamic capital allocation are quite limited in the literature. Bielecki et al. (2020) propose a backtesting method of ESCs based on the concept called fairness. This method is essentially treated in the more general framework of calibration and traditional backtests introduced in Nolde and Ziegel (2017). Despite the usefulness of such traditional backtests for model validation, Nolde and Ziegel (2017) also point out several issues of such tests particularly in terms of model comparison and banking regulation. To this end, our first aim in this paper is to present the comprehensive treatment of backtesting ESCs in the framework of traditional and comparative backtests (Nolde and Ziegel 2017) based on the concepts of joint identifiability and multi-objective elicitability (Fissler and Hoga 2024).
@@
Although the results shown in Fissler and Hoga (2024) are on systemic risk measures such as CoVaR, CoES and MES, in Section 2.2 we translate their results under the setting of capital allocation. Summarizing their results in our setting, ESCs are jointly identifiable with total VaR (i.e., VaR of the aggregated loss); moreover, ESCs themselves are not elicitable, but are multi-objective elicitable, when combined with total VaR and with respect to the lexicographical order. The last statement means more precisely that the tuple of true ESCs and total VaR is elicited as a unique minimizer of an expectation of some $\mathbb{R}^{2}$-valued scoring function with respect to the order on $\mathbb{R}^{2}$ such that the forecast evaluation of VaR is prioritized, and that of ESCs is conducted secondarily. Therefore, in the framework of a comparative backtest based on multi-objective elicitability it is necessary to forecast ESCs together with total VaR, and two forecasts of ESCs are comparable only when the corresponding forecasts of total VaR are equally accurate. Section 2.3 summarizes the framework of backtests of dynamic ESCs.
@@
When conducting a comparative backtest in practice, it may not always be clear which scoring function to use. In addition, the resulting forecast ranking can be sensitive to the choice of scoring functions (Patton 2020). To overcome these problems, a diagnostic tool called the Murphy diagram is explored in Ehm et al. (2016) and Ziegel et al. (2020) for robust forecast evaluation of VaR and ES, respectively, against the choice of scoring functions. Based on a mixture representation of a scoring function, the Murphy diagram is beneficial for checking whether one forecast dominates another under a class of scoring functions.
@@
As a second aim of this paper, we complement the framework of backtesting ESCs by developing their Murphy diagrams. They serve as a visual, intuitive way to compare competing forecasting models, thereby making it easier to digest and make informed decisions based on their performance. Section 3 introduces Murphy diagrams of ESCs and summarizes their properties.
@@
Our last aim of the paper is to leverage the concept of multi-objective elicitability and introduce a novel semiparametric model of dynamic ESCs. Section 4 proposes such a model, which combines the joint semiparametric estimation of the pair of total VaR and total ES with the compositional regression for modeling the dynamics of the proportions of total ES to each component of risk. Based on multi-objective elicitability of ESCs, forecast accuracy of total VaR is first evaluated prior to that of ESCs. In addition, risk allocation is conceptually a second stage in the risk management process after quantifying overall risk. This theoretical and practical two-stage procedure motivates us to first model the dynamics of the pair of total VaR and total ES and then estimate ESCs through the proportion of each source of risk to total ES. Since the component-wise sum of the vector of proportions must equal $1$, we model its dynamics by the compositional regression, which is a multiple regression for such compositional data.
@@
Note that Boonen et al. (2019) also apply compositional data analysis to the problem of capital allocation. They first estimate allocated capital from the losses under the assumption of normality and then fit and evaluate compositional regression models by regarding the set of normalized allocated capital as compositional data. One potential challenge of this procedure is that statistical error and model uncertainty reside not only in the compositional regression model, but also in the estimated allocated capital. To overcome this issue, we fit and evaluate compositional regression models based on a scoring function of ESCs. Multi-objective elicitability of ESCs allows us to quantify the accuracy of compositional regression models based on realized losses and not on estimated allocated capital. The major advantage of our proposed semiparametric model is in the complete separation of modeling ESCs from that of total VaR and total ES. This feature enables us to concentrate on modeling ESCs based on the existing model for the pair of total VaR and total ES and to assess the forecast accuracy of ESCs among the proposed compositional regression models and an existing one by equalizing their total VaR and total ES.
@@
Section 5 conducts an empirical analysis of a portfolio of stock returns and demonstrates superior performance of the proposed model compared with other models including those with conditional heteroskedastic volatilities, hysteretic effects, and time-varying correlations. Section 6 discusses potential directions for future research. We defer all technical results to Section S1, where references starting with the prefix “S” refer to the supplementary material. We also present details and additional results of our empirical analysis in Section S2.
@@ 2.1 Gradient allocations of expected shortfall
Throughout the paper we fix an atomless probability space $(\Omega,{\cal A},\mathbb{P})$, where all random objects are defined. For $p\in\{0\}\cup[1,\infty)$ and $d\in\mathbb{N}$, let ${\cal L}^{p}(\mathbb{R}^{d})$ be the set of all $\mathbb{R}^{d}$-value random vectors on $(\Omega,{\cal F},\mathbb{P})$ whose components have a finite $p$th moment. Let $F_{\bm{X}}$ be the joint cumulative distribution function (cdf) of $\bm{X}\in{\cal L}^{0}(\mathbb{R}^{d})$. We also denote by ${\cal F}^{p}(\mathbb{R}^{d})=\{F_{\bm{X}}:\bm{X}\in{\cal L}^{p}(\mathbb{R}^{d})\}$ and by ${\cal F}_{\text{c}}^{p}(\mathbb{R}^{d})$ the class of cdfs $F\in{\cal F}^{p}(\mathbb{R}^{d})$ with a strictly positive (Lebesgue) density for every $\bm{x}\in\mathbb{R}^{d}$ such that $F(\bm{x})\in(0,1)$. Analogously, ${\cal L}_{\text{c}}^{p}(\mathbb{R}^{d})$ denotes the set of $\mathbb{R}^{d}$-valued random vectors whose cdf belongs to ${\cal F}_{\text{c}}^{p}(\mathbb{R}^{d})$. For $X\in{\cal L}^{0}(\mathbb{R})$, VaR with a confidence level $\alpha\in(0,1)$ is given by $\operatorname{VaR}_{\alpha}(X)=\inf\{x\in\mathbb{R}:F_{X}(x)\geq\alpha\}$. Moreover, for $X\in{\cal L}^{1}(\mathbb{R})$, ES with a confidence level $\alpha\in(0,1)$ is $\operatorname{ES}_{\alpha}(X)=(1/(1-\alpha))\int_{\alpha}^{1}\operatorname{VaR}_{\beta}(X){\rm d}\beta$, which coincides with $\operatorname{ES}_{\alpha}(X)=\mathbb{E}[X|X\geq\operatorname{VaR}_{\alpha}(X)]$ if $X\in\mathcal{L}_{\text{c}}^{1}(\mathbb{R})$. The confidence level $\alpha$ is typically chosen to be close to $1$, such as $0.975$ and $0.99$. Let $\bm{X}=(X_{1},\dots,X_{d})^{\top}\in{\cal L}^{0}(\mathbb{R}^{d})$ be a random vector standing for the collection of losses to a portfolio or a financial institution. Moreover, let $S=X_{1}+\cdots+X_{d}$ be the total loss. Our sign convention is that positive values are losses and negative values are profits.
Under the prevalent Euler principle, the contribution of each component $X_{j}$, $j=1\dots,d$, to the risk of the total loss $\operatorname{ES}_{\alpha}(S)$ is determined by the gradient $(\partial/\partial\lambda_{j})\operatorname{ES}_{\alpha}(\bm{\lambda}^{\top}\bm{X})|_{\bm{\lambda}=\mathbf{1}_{d}}$ provided that the partial derivative exists, where $\bm{\lambda}=(\lambda_{1},\dots,\lambda_{d})^{\top}\in\mathbb{R}^{d}$ and $\mathbf{1}_{d}=(1,\dots,1)^{\top}\in\mathbb{R}^{d}$. Given certain smoothness conditions (Tasche 1999), this derivative leads to:
@@
which we call ESC for the $j$th risk. Note that ESC itself is well-defined for $(X_{j},S)\in{\cal L}^{1}(\mathbb{R}^{2})$ without smoothness assumptions. If $S\in\mathcal{L}_{\text{c}}^{1}(\mathbb{R})$, then the following full allocation property (FAP) holds:
@@ 2.2 Multi-objective elicitability and joint identifiability of ES contributions
We introduce several properties required for backtesting ESCs. Let ${\cal F}\subseteq{\cal F}^{0}(\mathbb{R}^{d})$, $k,\,m\in\mathbb{N}$, and $A\subseteq\mathbb{R}^{k}$. A function $\mathbf{S}:A\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ is called $\mathcal{F}$-integrable if, for every $F\in\mathcal{F}$ and $\bm{r}\in A$, every component of the function $\bm{x}\mapsto\mathbf{S}(\bm{r},\bm{x})$ is integrable with respect to $F$. A functional $\bm{\varrho}:{\cal F}\rightarrow A$ is called multi-objective elicitable on ${\cal F}$ with respect to a total order $\preceq$ on $\mathbb{R}^{m}$ if there exists an $\mathcal{F}$-integrable function $\mathbf{S}:A\times\mathbb{R}^{d}\rightarrow\mathbb{R}^{m}$ such that, for every $F\in{\cal F}$, $\bm{\varrho}(F)$ is the unique minimizer of $\bm{r}\mapsto\mathbb{E}[\mathbf{S}(\bm{r},\bm{X})]$, $\bm{X}\sim F$, over $A$. The function $\mathbf{S}$ is called a (strictly $\mathcal{F}$-consistent multi-objective) scoring function for $\bm{\varrho}$. We simply call $\bm{\varrho}$ elicitable when a scoring function can be taken with $m=1$.
@@
For $j\in\{1,\dots,d\}$, the $j$th ESC (1) coincides with the marginal expected shortfall (MES) of $(S,X_{j})^{\top}$ considered in Fissler and Hoga (2024) provided that $(S,X_{j})^{\top}\in\mathcal{F}_{\text{c}}^{1}(\mathbb{R}^{2})$. This observation immediately yields the following results on the multi-objective elicitability and joint identifiability of ESCs shown in Theorem 4.2 and Theorem S.3.1 of  Fissler and Hoga (2024), respectively. In the following, the lexicographic order $\leq_{\text{lex}}$ is adopted as a total order on $\mathbb{R}^{2}$. For every $(a_{1},b_{1}),(a_{2},b_{2})\in\mathbb{R}^{2}$, we write $(a_{1},b_{1})\leq_{\text{lex}}(a_{2},b_{2})$ if $a_{1}<a_{2}$ or if $(a_{1}=a_{2}\text{ and }b_{1}\leq b_{2})$. Moreover, we define:
@@
    $\mathbf{S}_{j}((m_{j},v),\bm{x})$ $=(\operatorname{S}^{\text{VaR}}(v,s),\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x}))^{\top},$
@@
    where $s=\sum_{i=1}^{d}x_{i}$, $h:\mathbb{R}\rightarrow\mathbb{R}$ is a strictly increasing function, and $\phi_{j}:\mathbb{R}\rightarrow\mathbb{R}$ is a strictly convex differentiable function with derivative $\phi_{j}^{\prime}$ such that $\mathbf{S}_{j}$ is $\mathcal{F}$-integrable.
@@
    $\mathbf{S}((\bm{m},v),\bm{x})=\left(\operatorname{S}^{\text{VaR}}(v,s),\sum_{j=1}^{d}\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x})\right)^{\top},$
@@
    $\mathbf{V}_{j}((m_{j},v),\bm{x})$ $=(\operatorname{V}^{\text{VaR}}(v,s),\operatorname{V}_{j}^{\text{ESC}}((m_{j},v),\bm{x}))^{\top},$
@@
    $\mathbf{V}((\bm{m},v),\bm{x})$ $=(\operatorname{V}^{\text{VaR}}(v,s),\operatorname{V}_{1}^{\text{ESC}}((m_{1},v),\bm{x}),\dots,\operatorname{V}_{d}^{\text{ESC}}((m_{d},v),\bm{x}))^{\top},$
@@
As a concrete example of the scoring functions in Proposition 1, the choice $h(s)=s$ in $\operatorname{S}^{\text{VaR}}$ yields the well-known pinball loss $\operatorname{S}^{\text{VaR}}(v,s)=(\mathbbm{1}{\{s\leq v\}}-\alpha)(v-s)$, and $\phi_{j}(x)=x^{2}$ in $\operatorname{S}_{j}^{\text{ESC}}$ leads to the squared loss $\operatorname{S}_{j}^{\text{ESC}}((m_{j},v),\bm{x})=\mathbbm{1}\{s>v\}(x_{j}-m_{j})^{2}$; see Nolde and Ziegel (2017) and Fissler and Hoga (2024) for more examples. Finally, we defer some technical discussion on the scoring and identification functions in Section S1.1.